# CaloClouds3: Ultra-Fast Geometry-Independent Highly-Granular Calorimeter Simulation

### v1

@@ 4.2 Performance of CaloClouds3 at random angles and fixed energies
Having established that CaloClouds3 can replicate photons travelling along the $z$ axis, the next assessment looks at photons all across a barrel segment to evaluate behaviour of CaloClouds3 across all angles. The following observables are computed at the reconstruction level, after the model has been integrated into the simulation pipeline.
@@
We now fix the energy of the incident particle in increments of $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ from $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ to $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. For legibility the histograms only present in full $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $50~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. All other energies are included in summary statistics plotting in figure 9 and plotted collectively in figure 17.
@@
Figure 7: Energy spectrum of CaloClouds3 (CC3) at random angles and fixed incident energies. For legibility, the plot only shows three of the ten photon incident energies tested; $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $50~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. Layout is the same as in figure 4.
@@
Figure 7 shows examples of the different energy distributions. The leftmost plot shows the per-cell energy distribution, with the far left blocked off by the MIP cut applied in reconstruction. For the majority of the spectrum, all three energies produce a flat ratio, only somewhat overestimating the tail of high energy cells.
The central plot of figure 7 shows the radial profile of the energy, and in the right hand plot the energy per layer is seen. At all energies, these distributions are well replicated, but the ratio is seen to improve somewhat at higher incident particle energies.
@@
Figure 8: Occupancy spectrum of CaloClouds3 (CC3) at random angles and fixed energies. For legibility, the plot only shows three of the ten photon incident energies tested; $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $50~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. Layout is the same as in figure 5.
@@
Figure 8, shows examples of the occupancy spectrum of CaloClouds3 at random angles and fixed energies. The leftmost plot is a histogram of total occupancy in each shower, while all three energies show reasonable agreement, it’s clear that the model performs better at higher energies.
@@
The central plot of figure 8 shows the radial profile of the occupancy, and in the right hand plot the occupancy per layer can be seen. At all energies, these distributions are well replicated, with a more pronounced deviation at $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ in the occupancy per layer.
@@
energy per
cell
energy per
layer
energy per
unit radius
occupancy
per layer
occupancy
per unit
radius
total
occupancy
@@ 4.3 Angular reconstruction
A common method for calculating the internal angle of the shower is to take the first component of an energy weighted PCA (principle component analysis) applied to all the reconstructed hits in the shower. When calculating the internal angle by PCA of hits, we observed systematic biases in the angle calculated in this way. This bias is caused by a combination of factors; undetected areas of secondary particles before the first layer of the detector, and the difference in path length of secondary particles raveling as shallow and steep angles. Further discussion, and illustration of these biases can be found in the appendix section C.
@@
Where $\vec{p}_{\mathrm{PCA}}$ is the direction of the photon reconstructed by PCA, and $\vec{p}_{\mathrm{inc}}$ is the true direction of the incident photon. Or alternatively, broken down into a $\phi$ component that rotates in the plane perpendicular the beam axis and a $\theta$ component that measures the angle to the beam axis;
@@
As the bias on this error is due to off-axis secondary hits, the simplest way to remove it was to do a PCA with only the highest energy hits in the shower. To rephrase this, rather than calculating the PCA on all hits in the shower, a limited percentage of the highest energy hits are selected, and a PCA is calculated on this subset. A scan over the percentage used is shown in figure 10.It is seen that not only does the error itself decrease, but the variance of the error is also reduced. With this we justify our preference for calculating the internal angle of the shower using the highest $4\%$ of hit energies in the shower.
@@
Figure 11: Angular errors of CaloClouds3 compared to Geant4. The left hand plot compares error in the azimuthal direction about the detector barrel, and the central plot compares error in the polar direction about the detector barrel (angle to the beam). The right hand plot compares the angular distance between the reconstructed internal angle and the true direction of the incident particle.
Despite the issues with modelling the angle of the lower energy hits, the agreement between CaloClouds3 and Geant4 is relatively good when calculating the internal angle using the classical PCA on $100\%$ of the hits as can be seen in the ratio. When using the improved metric for internal angle, the agreement with Geant4 becomes excellent. Summary statistics are shown in table 4. As the internal angle is a key feature in searches of long lived particles, replication of this metric provides additional use cases.
$\phi$
$\theta$
angular error
@@ 5 Di-Photon Separation
Reconstruction performance on a wider variety of metrics, across a range of fast simulation models, will be presented in a companion paper to this one. To validate this model with a basic example, we show the performance of CaloClouds3 on di-photon separation.
@@
Figure 13: The reconstructed multiplicity of two photons fired into the same barrel segment, as modelled in CaloClouds3 and Geant4. Three different energy are shown, from left to right: $5~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $20~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$.
In figure 13, we can see that CaloClouds3 reproduces the behaviour of Geant4 to within the fluctuation expected from the errors. In an analysis performed on samples generated by CaloClouds3 we have confidence that the photon multiplicity would be correctly reconstructed.
@@ 6 Inference Speed
The results are shown in Figure 14. This timing is done on a complied model and includes hit placement into the detector geometry.Unfortunately, it is not currently possible to utilise a GPU in this setup, but this matches the current realistic use pattern, in which fast simulation is mainly done on CPU.The CPU used for our tests was a AMD EPYC 7513 32-Core Processor from DESY’s Maxwell [58] cluster.
@@ Appendix A CaloClouds2 Architecture
In this appendix, we briefly describe the penultimate version, CaloClouds2555Previously written as CaloClouds II. A complete description, along with a description of the training data, is available in the existing paper [51].
@@
    The Center of Gravity (CoG) of the shower in the $x,y,z$ directions, rescaled to have a mean of 0 and standard deviation of $1$.
@@
This output is is used to calculate physical values in the Flow Calibration. The diffusion model is then conditioned on the incident particle energy and the total number of points predicted by ShowerFlow. The conditioned diffusion model creates a set of raw points, drawn from an iid666independent and identically distributed distribution. This diffusion model has been distilled such that it generates points in a single network evaluation. Distillation was not seen to cause any significant loss of accuracy as shown in [56].
The raw points created by the diffusion model are calibrated in the Final Calibration. Starting from the first detector layer the photon encounters and working outwards:
@@ Appendix C Angular error in PCA
In this section, we explore the structure of the error between the calculated internal angle and the true direction of the incident particle as described in section 4.3.
@@
Figure 18 shows the bias in the $\theta$ direction (polar angle) of the calculated internal angle, for both Geant4 and CaloClouds3. In the first column, where $100\%$ of the hits are used, the bias is clearly symmetric; in the positive $\theta$ direction, the error is positive, and in the negative $\theta$ direction, the error is negative. This is a physical result of the secondary particles encountering more active material at shallower angles, and clipping of the shower at the calorimeter surface, both of which rotate the first PCA component away from the true direction of the incident particle. This artefact is present in both Geant4 and CaloClouds3, but significantly more pronounced in CaloClouds3, hence it has a double peak in figure 11.
@@ 4 Acknowledgements
We thank Dirk Zerwas for the valuable comments on the manuscript. This research was supported in part by the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This project has received funding from the European Union’s Horizon 2020 Research and Innovation programme under Grant Agreement No 101004761. We acknowledge support by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and via the KISS consortium (05D23GU4, 13D22CH5) funded by the German Federal Ministry of Education and Research BMBF in the ErUM-Data action plan. A.K. has received support from the Helmholtz Initiative and Networking Fund’s initiative for refugees as a refugee of the war in Ukraine. P.M. has benefited from support by the CERN Strategic R&\&D Programme on Technologies for Future Experiments [59].

### v2

@@ 4.2 Performance of CaloClouds3 at random angles and fixed energies
Having established that CaloClouds3 can replicate photons travelling along the $z$ axis, the next assessment looks at photons all across a barrel segment to evaluate the behaviour of CaloClouds3 across all angles. The following observables are computed at the reconstruction level, after the model has been integrated into the simulation pipeline.
@@
We now fix the energy of the incident particle in increments of $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ from $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ to $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. For legibility the histograms only present in full $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $50~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. All other energies are included in summary statistics plotted in figure 9 and plotted collectively in figure 17.
@@
Figure 7: Energy spectrum of CaloClouds3 (CC3) at random angles and fixed incident energies. For legibility, the plot only shows three of the ten photon incident energies tested; $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $50~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. In the upper plot, Geant4 is shown as a solid line, and CaloClouds3 as a dashed line, though for some distributions the lines are too close to distinguish. The lower plot shows ratios between Geant4 and CaloClouds3. Layout is otherwise the same as in figure 4.
@@
Figure 7 shows examples of the different energy distributions. The leftmost plot shows the per-cell energy distribution, with the far left blocked off by the MIP cut applied in reconstruction. For the majority of the spectrum, all three energies produce a flat ratio, only somewhat overestimating the tail of high energy cells. The central plot of figure 7 shows the radial profile of the energy. Detector cells are assigned to a bin of this histogram according to their radial distance from the shower axis, and the energy from that cell added to the bin total. In the right hand plot the energy per layer is seen. At all energies, these distributions are well replicated, but the ratio is seen to improve somewhat at higher incident particle energies.
@@
Figure 8: Occupancy spectrum of CaloClouds3 (CC3) at random angles and fixed energies. For legibility, the plot only shows three of the ten photon incident energies tested; $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $50~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. In the upper plot, Geant4 is shown as a solid line, and CaloClouds3 as a dashed line, though for some distributions the lines are too close to distinguish. The lower plot shows ratios between Geant4 and CaloClouds3. Layout is otherwise the same as in figure 5.
@@
Figure 8, shows examples of the occupancy spectrum of CaloClouds3 at random angles and fixed energies. The leftmost plot is a histogram of total occupancy in each shower, while the higher two energies show reasonable agreement, it’s clear that the model does not perform as well at the lowest energy.
@@
The central plot of figure 8 shows the radial profile of the occupancy. Like the radial energy, detector cells are assigned to a bin of this histogram according to their radial distance from the shower axis, and the number of cells with energy above half a MIP is counted towards the bin total. Towards 0 radius, bins encompass a smaller volume of the detector, creating a dip in the number of active cells. In the right hand plot the occupancy per layer can be seen. At all energies, these distributions are well replicated, with a more pronounced deviation at $10~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ in the occupancy per layer.
@@
energy per cell energy per layer energy per unit radius occupancy per layer occupancy per unit radius total occupancy
@@ 4.3 Angular reconstruction
A common method for calculating the internal angle of the shower is to take the first component of an energy weighted PCA (principle component analysis) applied to all the reconstructed hits in the shower. When calculating the internal angle by PCA of hits, we observed systematic biases in the angle calculated in this way. This bias is caused by a combination of factors. Firstly, the recorded shower misses any secondary particles that were emitted before the first layer of the detector. When the shower is at an angle, these missed particles are no longer symmetrically distributed about the shower axis. Secondly, secondary particles travelling at steep and shallow angles experience different path lengths in the detector. When the shower axis is at an angle, the secondary particles from one side of the axis experience a different path length to those emitted on the other side, causing an imbalance in their detection. Further discussion, and illustration of these biases can be found in the appendix section C.
@@
Where $\vec{p}_{\mathrm{PCA}}$ is the direction of the photon reconstructed by PCA, and $\vec{p}_{\mathrm{inc}}$ is the true direction of the incident photon. Or alternatively, broken down into a $\phi$ component that rotates in the plane perpendicular to the beam axis and a $\theta$ component that measures the angle to the beam axis;
@@
As the bias on this error is due to off-axis secondary hits, the simplest way to remove it is to do a PCA with only the highest energy hits in the shower. To rephrase this, rather than calculating the PCA on all hits in the shower, a limited percentage of the highest energy hits are selected, and a PCA is calculated on this subset. A scan over the percentage used is shown in figure 10. It is seen that not only does the error itself decrease, but the variance of the error is also reduced. With this we justify our preference for calculating the internal angle of the shower using the highest $4\%$ of hit energies in the shower. Metrics calculated with only this highest energy fraction are referred to as improved metrics.
@@
Figure 11: Angular errors of CaloClouds3 compared to Geant4. The left hand plot compares error in the azimuthal direction about the detector barrel, and the central plot compares error in the polar direction about the detector barrel (angle to the beam). The right hand plot compares the angular distance between the reconstructed internal angle and the true direction of the incident particle. Each plot, shows showers from two models, and two calculations are given, one using all of the hits, and then the improved metric, using only the highest energy $4\%$ of the hits.
Despite the issues with modelling the angle of the lower energy hits, the angular distance distributions of CaloClouds3 and Geant4 remain within a factor of $2$ when calculating the internal angle using the classical PCA on $100\%$ of the hits as can be seen in the ratio. When using the improved metric for internal angle, the agreement with Geant4 becomes good for all angular metrics, including a proper replication of the $\theta$ shape. Summary statistics are shown in table 4; every single metric improves at $4\%$. Due to the shape of the barrel segments, the sample spans a smaller range in $\phi$, compared to $\theta$, see section 2.2, therefore it is to be expected that the $\theta$ and overall angular error metrics will improve more than $\phi$. As the internal angle is a key feature in searches of long lived particles, replication of this metric provides additional use cases.
$\phi$ $\theta$ angular error
@@ 5 Di-Photon Separation
Reconstruction performance on a wider variety of metrics, across a range of fast simulation models, is presented in a companion paper to this one, a first full physics benchmark for highly granular calorimeter surrogates [16]. To validate this model with a basic example, we show the performance of CaloClouds3 on di-photon separation.
@@
Figure 13: The reconstructed multiplicity of two photons fired into the same barrel segment, as modelled in CaloClouds3 and Geant4. Three different energies are shown, from left to right: $5~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, $20~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $100~$\mathrm{G}\mathrm{e}\mathrm{V}$\$. The shaded grey and orange bands in the top plot represent the errors of Geant4 and CaloClouds3 respectively, while the shaded orange band in the ratio plot below represents the propagated combined error on the ratio.
In figure 13, we can see that CaloClouds3 reproduces the behaviour of Geant4 to within the fluctuations expected from the errors. In an analysis performed on samples generated by CaloClouds3 we have confidence that the photon multiplicity would be correctly reconstructed.
@@ 6 Inference Speed
The results are shown in Figure 14. This timing is done on a compiled model and includes hit placement into the detector geometry. Unfortunately, it is not currently possible to utilise a GPU in this setup, but this matches the current realistic use pattern, in which fast simulation is mainly done on CPU. The CPU used for our tests was a AMD EPYC 7513 32-Core Processor from DESY’s Maxwell [22] cluster.
@@ Appendix A CaloClouds2 Architecture
In this appendix, we briefly describe the penultimate version, CaloClouds2666Previously written as CaloClouds II. A complete description, along with a description of the training data, is available in the existing paper [11].
@@
    The CoG of the shower in the $x,y,z$ directions, rescaled to have a mean of 0 and standard deviation of $1$.
@@
This output is is used to calculate physical values to Apply macro features. The diffusion model is then conditioned on the incident particle energy and the total number of points predicted by ShowerFlow. The conditioned diffusion model creates a set of raw points, drawn from an iid distribution. This diffusion model has been distilled such that it generates points in a single network evaluation. Distillation was not seen to cause any significant loss of accuracy as shown in [15].
The raw points created by the diffusion model are rescaled to Apply macro features. Starting from the first detector layer the photon encounters and working outwards:
@@ Appendix C Angular error in PCA
In this section, we explore the structure of the error between the calculated internal angle and the true direction of the incident particle as described in section 4.3. For each shower, an internal direction is calculated using PCA. Either the calculated uses all the hits, or for the improved metric, only the $4\%$ highest energy hits are used. This can then be compared to the true direction of the incident particle, and the difference between the calculate and true direction is refered to as a bias. For example, if the PCA direction has lower $\theta$ than that of the incident particle, the bias in $\theta$ will be negative.
@@
Figure 18 shows the bias in the $\theta$ direction (polar angle) of the calculated internal angle, for both Geant4 and CaloClouds3. In the first column, where $100\%$ of the hits are used, the bias is clearly symmetric; in the positive $\theta$ direction, the error is positive, and in the negative $\theta$ direction, the error is negative. This is a physical result of the secondary particles encountering more active material at shallower angles, and clipping of the shower at the calorimeter surface, both of which rotate the first PCA component away from the true direction of the incident particle. This artifact is present in both Geant4 and CaloClouds3, but significantly more pronounced in CaloClouds3, hence it has a double peak in figure 11.
@@ 4 Acknowledgements
We thank Dirk Zerwas for the valuable comments on the manuscript. This research was supported in part by the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This project has received funding from the European Union’s Horizon 2020 Research and Innovation programme under Grant Agreement No 101004761. We acknowledge support by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and via the KISS consortium (05D23GU4, 13D22CH5) funded by the German Federal Ministry of Education and Research BMBF in the ErUM-Data action plan. A.K. has received support from the Helmholtz Initiative and Networking Fund’s initiative for refugees as a refugee of the war in Ukraine. P.M. has benefited from support by the CERN Strategic R&\&D Programme on Technologies for Future Experiments [36]. Finally, we thank the anonymous referees from JINST for their through and thoughtful reading, which was a great assistance in polishing this work.